ElasticHosts Help Centre

packet loss in Toronto

Resolved May 15th, 13:55 - May 16th, 10:00 BST

Timeline

Updateabout 9 days ago

The issue has been resolved. We are awaiting the RFO from the datacenter. If you would like a copy, please open a support ticket.

Monitoringabout 10 days ago

There is an ongoing fibre cut issue.

Engineers have stabilised the network by re-routing traffic via our redundant links so all packet loss should've stopped.

We'll update this status when normal routing has resumed.

Monitoringabout 10 days ago

Network connectivity is back to normal now. We are still investigating the cause.

Investigatingabout 10 days ago

We are seeing up to 50% packet loss in Toronto and are currently investigating the cause